# Efficient Multi-agent Navigation with Lightweight DRL Policy

### v1

@@ Abstract
In this article, we present an end-to-end collision avoidance policy based on deep reinforcement learning (DRL) for multi-agent systems, demonstrating encouraging outcomes in real-world applications. In particular, our policy calculates the control commands of the agent based on the raw LiDAR observation. In addition, the number of parameters of the proposed basic model is 140,000, and the size of the parameter file is 3.5 MB, which allows the robot to calculate the actions from the CPU alone. We propose a multi-agent training platform based on a physics-based simulator to further bridge the gap between simulation and the real world. The policy is trained on a policy-gradients-based RL algorithm in a dense and messy training environment. A novel reward function is introduced to address the issue of agents choosing suboptimal actions in some common scenarios. Although the data used for training is exclusively from the simulation platform, the policy can be successfully transferred and deployed in real-world robots. Finally, our policy effectively responds to intentional obstructions and avoids collisions. The website is available at https://sites.google.com/view/xingrong2024efficient/%E9%A6%96%E9%A1%B5.
keywords:
@@
Multi-agent systems, distributed collision avoidance, reactive and sensor-based planning, simulation to real gap.
1 Introduction
@@
Multi-agent systems are increasingly used in robotics and artificial intelligence, such as industrial automation (e.g., ), material transportation (e.g., Xie et al. (2024)), and agricultural automation (e.g., Dutta et al. (2021)). The main challenge is implementing an efficient, robust, and safe navigation algorithm for multiple robots so that the robot can safely and efficiently complete point-to-point movement.
@@
The traditional approaches for solving the problem of multi-agent navigation are primarily categorized into two types. One of the approaches is the centralized approach. These approaches calculate trajectories for each agent through a central server using the state of all agents and their respective workspaces (e.g., Tang et al. (2018), Zhou et al. (2021)). However, the computational cost associated with these approaches escalates as the number of agents increases, along with the maintenance of a more extensive communication network. In addition, once the central server is attacked, all agents will stop working. The other approach is the decentralized approach, which is deployed to each agent, empowering them with the capability to make decisions (e.g., Reciprocal Velocity Obstacle (RVO, Fiorini and Shiller (1998)), Optimal Reciprocal Collision Avoidance (ORCA, van den Berg et al. (2011))). However, these algorithms rely on a perfect perception of the environment and the state of the other agents, posing challenges for their application in larger-scale scenarios and deployment onto real-world robots. Numerous variants have been proposed to address these limitations (e.g., Alonso-Mora et al. (2010), Godoy et al. (2016)). Additionally, RVO and its variants often necessitate tuning numerous hyperparameters to achieve optimal performance, and the adjustment of these parameters is highly dependent on the specific scenario.
@@
Learning-based approaches have been proposed for robot navigation in recent years. Some approaches use imitation learning to learn from expert data and work successfully (e.g., Karnan et al. (2022), Yan et al. (2023)). However, the effectiveness of imitation learning depends on expert data quality, and it is challenging to handle scenarios out of the expert data. Other approaches employ reinforcement learning (RL) or deep reinforcement learning (DRL) to train policy. Some RL-based approaches achieve desired outcomes by modifying the reward function (e.g., Xie and Dames (2023), Han et al. (2022)). These approaches can yield great results in corresponding scenarios, but some suffer from overly redundant reward function designs and struggle to handle unforeseen scenarios. Some approaches train agents in specific scenarios to learn more diverse policies (e.g., Fan et al. (2020), Xue and Chen (2024)). Learning in more specific scenarios can enrich policies but demands higher standards for scenario design. Besides, there are efforts to bridge the gap between simulation and the real world by using progressive networks or adding randomness to the simulator (e.g., Rusu et al. (2016), Xie et al. (2023)). Such approaches enable better deployment of policies learned from simulators on real robots, yet they do not fully leverage physics-based simulators. Finally, as the robot in a decentralized system can only perceive local observations during navigation, the decentralized robot navigation can be defined as the partially observable Markov decision process (POMDP). Most policies use convolutional neural networks (CNN) to encode perceptual information (e.g., Liu et al. (2020), Chen et al. (2020)), which results in the loss of historical information during navigation. Some approaches propose using Long Short-Term Memory (LSTM, Hochreiter and Schmidhuber (1997)) to record memory, providing more comprehensive information for navigation (e.g., Liu et al. (2021), Cui et al. (2022)).
In this paper, we propose a lightweight end-to-end distributed multi-agent navigation policy that maps raw LiDAR observation into action commands for the agents. We use the LiDAR sensor in perception for the following reasons (Zhu and Zhang (2021)): 1. the gap of the LiDAR sensor between the simulation and the real world is close; 2. the small measurement dimension of LiDAR is beneficial for simulation-to-real transfer. Furthermore, we introduce a physics-based simulation platform tailored for multi-agent training to bridge the gap between simulation and the real world. We have successfully deployed our policy in simulation and real-world experiments and tested the policy’s performance in single-agent and multi-agent scenarios. Additionally, the policy can avoid intentional obstructions, like a rapid-moving obstacle controlled by a human.
@@
  • We propose a multi-agent training platform based on a three-dimensional (3-D) simulator with a physical engine to bridge the gap between simulation and the real world. We compare a two-dimensional (2-D) simulator without a physics engine and a 3-D simulator with a physics engine with real-world environments. The result proves that a 3-D simulator with a physics engine can provide better randomness for agent training.
  • We propose a lightweight end-to-end distributed multi-agent navigation policy, which utilizes the hidden state in the LSTM to alleviate the information loss in the POMDP. Besides, an attention structure based on the lightweight policy is proposed to improve the policy’s performance in multi-agent scenarios. Both policies are trained in a stochastic environment and can be directly deployed to real-world robots.
  • We propose a novel reward function to solve the problem: the agent may choose suboptimal actions in some common scenarios, such as passing through narrow gaps and multi-agent travels in the opposite direction.
@@
The rest of this article is organized as follows. Section 2 is a brief review of the related work. Section 3 introduces the notation and formulation used in this article. Section 4 discusses the policy’s structure, the RL setup for training multi-agent navigation policy, and the simulator platform. Section 5 presents the policy’s performance in the simulator and real-world experiments. Finally, we conclude our work and propose directions for future improvement in Section 6.
2 Related work
@@
In this section, we briefly describe some related works of multi-agent navigation. These works can be categorized into trajectory-based, reaction-based, and learning-based approaches. The learning-based approaches can be further divided into supervised learning and RL. Finally, we discuss the approaches to bridge the gap between simulation and the real world.
2.1 Trajectory-based approaches
@@
Trajectory-based approaches generally require detailed environmental information to plan a feasible trajectory for the robot (e.g., Model Predictive Control (MPC, Keviczky et al. (2004)) and Dynamic Window Approach (DWA, Fox et al. (1997))). In recent years, new theories, such as safe corridors, buffered Voronoi cells (BVC), and Graphs of Convex Sets (GCS), have been introduced to transform the navigation problem into a convex optimization problem. Park et al. (2023) utilizes a dynamic linear safe corridor while introducing mode-based subgoal planning to tackle planning problems in complex environments. They also use collision alert propagation and escape point planning to deal with emergencies. Zhu et al. (2022) constructs a chance-constrained buffered uncertainty-aware Voronoi cell (B-UAVC) for each robot and ensures the safety of robot navigation by specifying thresholds. Marcucci et al. (2023) proposes a framework that enables convex optimization to effectively and reliably plan trajectories around obstacles through a practical convex relaxation programming problem by using GCS. Although the trajectory-based algorithm can obtain the optimal trajectory, it will lead to large computational costs or difficult convex optimization when applied in large or unstructured scenarios.
2.2 Reaction-based approaches
@@
Reaction-based approaches accomplish navigation tasks by continuously sensing and responding to the surrounding environment, enabling them to adapt to dynamically changing conditions. Reciprocal velocity obstacles (RVO) framework (Fiorini and Shiller (1998)) maps the dynamic environment into the robot velocity space and generates safe control velocities from these velocity constraints. The ORCA framework (van den Berg et al. (2011)) follows the principles of RVO while resolving the jitter problem in RVO and improves the obstacle avoidance efficiency. Moreover, Alonso-Mora et al. (2010) further guarantees smooth, collision-free motions under non-holonomic constraints. However, these approaches struggle to be implemented in the real world because they require a perfect perception of the agent’s surroundings. In response to these limitations, some variants have been developed. Godoy et al. (2016) enhances environmental awareness through communication among agents, allowing for more accurate and timely adjustments based on shared information. Arul and Manocha (2021) incorporates Buffered Voronoi Cells to implement a more conservative obstacle avoidance policy, ensuring a safe distance from obstacles and other agents. To some extent, these variants have solved the perception problem. Still, ORCA and its variants are sensitive to parameters, and the corresponding parameters are different in different scenarios, which makes these approaches difficult to adjust in practical applications.
2.3 Supervised learning-based approaches
With the advancement of machine learning, numerous studies utilize it to solve the robot navigation problem. Supervised learning-based approaches train the policy with extensive datasets, either through experts or traditional algorithms, and enable agents to learn how to map observations into actions. Yan et al. (2023) adopts an imitation learning approach, training a mapless navigation policy solely from 2-D LiDAR inputs, incorporating an aggregated dataset with additional safety enhancements. Karnan et al. (2022) utilizes imitation learning to devise a purely vision-based navigation algorithm, successfully learning navigation policies from a single video demonstration collected by physically distinct agents. Cho and Chung (2024) uses a risk-sensitive approach to prioritize learning on scarce but high-risk terrain through self-supervised learning and enriches reliable data for inexperienced terrain with pseudo-labels. Though supervised learning-based approaches have achieved significant breakthroughs in robot navigation, they struggle to cope with scenarios not represented in the training dataset. Furthermore, the reliance on expert knowledge inherently limits the policies from surpassing the boundaries of that preexisting knowledge.
2.4 Reinforcement learning-based approaches
@@
Generally, the reward function of the navigation problem is similar to the potential field function. Chen et al. (2017b) proposes a dense reward function that can guide reinforcement learning concisely and effectively. However, the reward function will cause the agent to learn suboptimal actions, which we will discuss in Section 4. Some approaches define the reward function through other rules. Chen et al. (2017a) develops a time-efficient navigation policy that respects public social norms by inducing the right-handed rules in the reward function. Xie and Dames (2023) introduces a velocity obstacle term into the reward function, enabling mobile robots to navigate autonomously in spaces filled with static obstacles and dense pedestrian traffic.
Some approaches train the policy in specific international scenarios. Li et al. (2019) proposes an advanced version of the Social Attention Reinforcement Learning (SARL) algorithm, achieving human-aware navigation in indoor environments. Yao et al. (2020) employs a grid map-based sensor fusion approach to map three frames of egocentric grid maps and the robot’s relative local target positions into low-level robotic control commands. Huang et al. (2024) provides the agent with a curriculum and a replay buffer of clipped collision segments, enhancing performance in obstacle-rich environments.
@@
Regarding the neural network structure, some policies use CNNs to encode perceptual information, such as Liu et al. (2020), Guldenring et al. (2020), Chen et al. (2021). To enhance the policy’s utilization of past states, some approaches harness the hidden state structure of RNNs to store historical information. Liu et al. (2021) presents the Decentralized Structured Recurrent Neural Network (DS-RNN), which uses spatial and temporal relationships for robotic decision-making in swarm navigation. Xue and Chen (2024) introduces a novel multi-agent recurrent deterministic policy gradient algorithm and leverages LSTM to improve action value predictions using historical exploration information.
@@
Some traditional control approaches have also been employed to enhance the safety of RL training. Cheng et al. (2019) utilizes a model-based controller leveraging Control Barrier Functions (CBFs) and online learning of unknown system dynamics to ensure safety during the learning phase. Emam et al. (2022) incorporates safety as a differentiable robust control barrier function layer within a model-based RL framework and proposes a modular approach to learning potential reward-driven tasks independently of safety constraints.
2.5 Simulation-to-real gap
@@
As training is inefficient in the real world and easily damages the robot, RL-based approaches are usually trained in simulators. Kahn et al. (2021) proposed an approach to collect data in reality and train policy that can navigate in real-world urban and off-road environments with geometrically distracting obstacles. However, the gap between simulations and real-world environments remains a crucial barrier preventing the widespread adoption of RL-based approaches in real robots. Several approaches have been proposed to address this issue and bridge the simulation-to-reality gap. Rusu et al. (2016) utilizes progressive networks to bridge the reality gap and transfer policies learned in simulation to the real world. Tobin et al. (2017) investigates domain randomization, a straightforward technique for training models on simulated images that transfers to real images by randomizing renderings within the simulator. García and Molina (2019) proposes an approach to evaluate LiDAR sensors and navigation and obstacle avoidance capabilities in simulated scenarios using a real drone platform. Fan et al. (2020) proposes a hybrid control RL approach, which initially learns an end-to-end general obstacle avoidance policy through RL. Subsequently, hybrid control is incorporated into this policy to improve its feasibility and robustness in real-world environments. This approach leverages the strengths of both RL for learning complex behaviors and hybrid control to ensure reliability and stability in practical applications. Shi et al. (2023) proposed a domain adaptation approach based on a cyclic consistent generative adversarial network and a force control transfer approach to bridge the reality gap.
@@
Our approach further bridges the gap between simulation and the real-world environment by designing a training platform based on a physics-based simulator. We also propose a novel policy network based on LSTM and attention mechanism, which is state of the art (SOTA) of the navigation based on pure LiDAR sensor. Furthermore, a novel reward function is designed to solve some common scenarios, such as agent passes through narrow gaps.
3 Problem formulation
@@
In this section, we will introduce the notation of this work and our reward function.
3.1 Notation of reinforcement learning

### v3

@@ Abstract
This article presents an end-to-end collision avoidance policy based on deep reinforcement learning (DRL) for multi-agent systems, demonstrating encouraging outcomes in real-world applications. In particular, our policy calculates the control commands of the agent based on the raw LiDAR observations. In addition, the number of parameters of the proposed basic model is 140,000, and the size of the parameter file is 3.5 MB, which allows the robot to calculate the actions from the CPU alone. We propose a multi-agent training platform based on a physics-enhanced simulator to further bridge the gap between simulation and the real world. The policy is trained on a policy-gradients-based RL algorithm in a dense and messy training environment. A novel reward function is introduced to address the issue of agents choosing suboptimal actions in some common scenarios. Although the data used for training is exclusively from the simulation platform, the policy can be successfully transferred and deployed in real-world robots. Finally, our policy effectively responds to intentional obstructions and avoids collisions. The website is aviliable in https://sites.google.com/view/xingrong2024efficient/%E9%A6%96%E9%A1%B5
Note to Practitioners
We’ve developed a navigation system that enables robots to navigate collision-free in dynamic environments such as warehouses, factories, and hospitals. It leverages LiDAR data to make rapid decisions with minimal computational resources, making it highly efficient for small robots. This system is trained in physics-enhanced simulations, ensuring its suitability for real-world applications and exceptional performance in handling challenging scenarios like tight spaces and unexpected obstacles. Whether you need robots to move efficiently in a bustling warehouse, transport materials in a factory, or assist in a hospital, this approach provides a practical solution for deploying effective navigation without requiring high-end hardware or intricate configurations.
Index Terms:
@@
Deep reinforcement learning, distributed collision avoidance, reactive and sensor-based planning, simulation to real gap.
I INTRODUCTION
@@
Multi-agent systems are increasingly used in robotics and artificial intelligence, such as industrial automation [1], material transportation [2]. The main challenge is implementing an efficient, robust, and safe navigation algorithm for multiple robots so that the robot can safely and efficiently complete point-to-point movement.
@@
The traditional approaches for solving the problem of multi-agent navigation are primarily categorized into two types. One of the approaches is the centralized approach. These approaches calculate trajectories for each agent through a central server using the state of all agents and their respective workspaces [3]. However, the computational cost and complexity of these approaches increase significantly with the number of agents, and the maintenance of a more extensive communication network. In addition, once the central server is attacked, all agents will stop working. The other approach is the decentralized approach, which is deployed to each agent, empowering them with the capability to make decisions, such as Reciprocal Velocity Obstacle (RVO) [4] and Optimal Reciprocal Collision Avoidance (ORCA) [5]. However, these algorithms rely on an idealized perception of the environment and the states of the other agents, posing challenges for their application in larger-scale scenarios and deployment onto real-world robots. Numerous variants have been proposed to address these limitations [6], [7]. Additionally, RVO and its variants often necessitate tuning numerous hyperparameters to achieve optimal performance, and the adjustment of these parameters is highly dependent on the specific scenario.
@@
Learning-based approaches have been proposed for robot motion planning in recent years [8]. Some approaches use imitation learning to learn from expert data and work successfully [9]. However, the effectiveness of imitation learning depends on expert data quality, and it is challenging to handle scenarios out of the expert data. Other approaches employ reinforcement learning (RL) or deep reinforcement learning (DRL) to train policy. Some RL-based approaches achieve desired outcomes by modifying the reward function [10], [11]. These approaches can yield excellent results in corresponding scenarios, but some suffer from redundant reward function designs and struggle to handle unforeseen scenarios. Some approaches train agents in specific scenarios to learn more diverse policies [12], [13]. Learning in more specific scenarios can enrich policies but demands higher standards for scenario design. Besides, there are efforts to bridge the gap between simulation and the real world by using progressive networks or adding randomness to the simulator [14], [15]. Such approaches enable better deployment of policies learned from simulators on real robots, yet they do not fully leverage physics-enhanced simulators. Finally, as the robot in a decentralized system can only perceive local observations during navigation, the decentralized robot navigation can be defined as the partially observable Markov decision process (POMDP). Most policies use convolutional neural networks (CNN) to encode perceptual information [16], [17], which results in the loss of historical information during navigation. Some approaches propose using long short-term memory (LSTM) [18] to record memory, providing more comprehensive information for navigation [19], [20].
This paper presents a lightweight, end-to-end distributed multi-agent navigation framework that directly maps raw LiDAR observations to agent action commands. LiDAR is selected as the primary perception modality due to its minimal simulation-to-reality gap and compact measurement dimensions, which facilitate effective simulation-to-real transfer [21]. We develop a physics-enhanced simulation platform replicating real-world dynamics to bridge the reality gap in multi-agent training. Extensive experimental validation demonstrates the policy’s robustness in simulated and physical environments, including single-agent and multi-agent scenarios. Notably, the framework successfully handles adversarial conditions, such as rapidly moving human-controlled obstacles.
@@
  • We develop a physics-enhanced simulation platform that bridges the simulation-to-reality gap. Comparisons show the physics-integrated system improves environmental randomness and training effectiveness.
  • We propose a lightweight distributed framework using LSTM hidden states to address POMDP information loss. An integrated attention mechanism enhances coordination, with policies trained in stochastic environments and transferable to real-world robots.
  • We introduce a reward function addressing classical formulation limitations, resolving suboptimal behaviors in narrow passages and bidirectional multi-agent navigation.
@@
The rest of this article is organized as follows. Section II is a brief review of the related work. Section III introduces the notation and formulation used in this article, the proposed reward function, the policy’s structure, the overall framework of the approach, and the physics-enhanced simulator platform. Section IV presents the policy’s performance in the simulation and real-world experiments. Finally, we conclude our work and propose directions for future improvement in Section V.
II Related Work
@@
In this section, we briefly describe some related works of multi-agent navigation. These works can be categorized into trajectory-based, reaction-based, and learning-based approaches. Finally, we discuss the approaches to bridge the gap between simulation and the real world.
II-A Trajectory-based Approaches
@@
Trajectory-based approaches generally require detailed environmental information to plan a feasible trajectory for the robot [22]. In recent years, new theories, such as safe corridors, buffered Voronoi cells (BVC), and Graphs of Convex Sets (GCS), have been introduced to transform the navigation problem into a convex optimization problem. Park [23] utilizes a dynamic linear safe corridor while introducing mode-based subgoal planning to tackle planning problems in complex environments. Zhu [24] constructs a chance-constrained buffered uncertainty-aware Voronoi cell (B-UAVC) for each robot and ensures the safety of robot navigation by specifying thresholds. Marcucci [25] proposes a framework that enables convex optimization to effectively and reliably plan trajectories around obstacles through a practical convex relaxation programming problem by using GCS. Although the trajectory-based algorithm can obtain the optimal trajectory, it will lead to significant computational costs or conplex convex optimization when applied in large or unstructured scenarios.
II-B Reaction-based Approaches
@@
Reaction-based approaches accomplish navigation tasks by continuously sensing and responding to the surrounding environment, enabling them to adapt to dynamically changing conditions. Reciprocal velocity obstacles (RVO) [4] maps the dynamic environment into the robot velocity space and generates safe control velocities from these velocity constraints. The ORCA [5] follows the principles of RVO while resolving the jitter problem in RVO and improves obstacle avoidance efficiency. Moreover, ORCA-NH [6] further guarantees smooth, collision-free motions under non-holonomic constraints. However, these approaches struggle to be implemented in the real world because they require a perfect perception of the agent’s surroundings. Still, ORCA and its variants are sensitive to parameters, and the corresponding parameters are different in different scenarios, which makes these approaches challenging to adjust in practical applications.
II-C Reinforcement Learning-based Approaches
@@
Generally, the reward function of the navigation problem is similar to the potential field function. Chen [26] proposes a dense reward function that can guide reinforcement learning concisely and effectively. Some approaches define the reward function through other rules. Xie [10] introduces a velocity obstacle term into the reward function, enabling mobile robots to navigate autonomously in spaces filled with static obstacles and dense pedestrian traffic. Some approaches train the policy in specific international scenarios.
@@
Regarding the neural network structure, some policies use CNNs to encode perceptual information [16]. To improve the policy’s ability to leverage historical states, some approaches utilize the hidden state architecture of RNNs to encode and store past information. Liu [19] presents the Decentralized Structured Recurrent Neural Network (DS-RNN), which uses spatial and temporal relationships for robotic decision-making in swarm navigation. Xue [13] introduces a novel multi-agent recurrent deterministic policy gradient algorithm and leverages LSTM to improve action value predictions using historical exploration information.
@@
Some traditional control approaches have also been employed to enhance the safety of RL training. Cheng [27] utilizes a model-based controller leveraging Control Barrier Functions (CBFs) and online learning of unknown system dynamics to ensure safety during the learning phase.
II-D Simulation-to-Real Gap
@@
As training is inefficient in the real world and easily damages the robot, RL-based approaches are usually trained in simulators. However, the gap between simulations and real-world environments remains a crucial barrier preventing the widespread adoption of RL-based approaches in real robots. Several methods have been proposed to address this issue and bridge the simulation-to-reality gap. Carcia [28] proposes an approach to evaluate Lidar sensors and navigation and obstacle avoidance capabilities in simulated scenarios using a real drone platform. Fan [12] proposes a hybrid control RL approach, which initially learns an end-to-end general obstacle avoidance policy through RL. Subsequently, hybrid control is incorporated into this policy to improve its feasibility and robustness in real-world environments.
@@
Our approach further bridges the gap between simulation and the real-world environment by designing a training platform based on a physics-based simulator. We propose a novel policy network based on LSTM and attention mechanism, a state-of-the-art (SOTA) navigation based on a pure LiDAR sensor. Furthermore, a novel reward function is designed to solve common scenarios, such as an agent passing through narrow gaps.
III Method
@@
In this section, we will introduce the notation of this work, the proposed reward function, the architecture of the policy network, the training process, and the simulation platform with training scenarios.
III-A Notations and Problem Statement